# Optimal transport for types and convex analysis for definable predicates in tracial $\mathrm{W}^{*}$-algebras

### 02 Jan 2024 v2

@@ 3.1. Definition and Properties
Thus, by Proposition 2.21, $\tilde{\mathcal{P}}\preceq\mathcal{M}$. Hence, $\tilde{\mathcal{P}}$ is an elementary submodel of $\mathcal{M}$ that contains $A$ but not $x$, which completes the proof. ∎
@@
because $\mathcal{M}\preceq\mathcal{N}_{1}$. Therefore, $\Phi$ is approximately satisfiable in $\mathcal{M}$, and hence $\Phi$ is satisfiable in $\mathcal{M}$ because $\mathcal{M}$ is $\kappa$-saturated. If $x\in D^{\mathcal{M}}$ satisfies $\Phi$, then $\operatorname{tp}^{\mathcal{N}_{1}}(x)=\operatorname{tp}^{\mathcal{M}}(x)=\operatorname{tp}^{\mathcal{N}_{1}}(y)$. Since $\mathcal{N}_{1}$ is strongly $\kappa$-homogeneous, there exists an automorphism $\Theta$ of $\mathcal{N}_{1}$ with $\Theta(x)=y$. By (1) and Proposition 3.3, $x\in\operatorname{acl}^{\mathcal{N}_{1}}(\varnothing)$, which is the intersection of all elementary submodels of $\mathcal{N}_{1}$ by Proposition 3.5. Therefore, because the automorphic image of an elementary submodel is also an elementary submodel, $y=\Theta(x)$ is in the intersection of all elementary submodels of $\mathcal{N}_{1}$, which is $\operatorname{acl}^{\mathcal{N}_{1}}(\varnothing)=\operatorname{acl}^{\mathcal{M}}(\varnothing)=\mathcal{M}$. Therefore, $\mathcal{N}_{1}=\mathcal{M}$ and also $\mathcal{N}=\mathcal{M}$. (3) Let $\mathcal{N}\preceq\mathcal{M}$ be a $\kappa$-saturated and strongly $\kappa$-homogeneous elementary extension. Then $\mathcal{M}=\mathcal{N}$, so we are done. ∎
@@ 3.2. Examples in von Neumann algebras
Definable and algebraic closures have not yet been studied in von Neumann algebras (except for their relationship with $1$-bounded entropy in [39, §4.5]), and therefore, we develop several basic properties and examples here. Given that the definable closure is always a $\mathrm{W}^{*}$-subalgebra by Corollary 3.6, we have $\operatorname{dcl}^{\mathcal{M}}(A)=\operatorname{dcl}^{\mathcal{M}}(\mathrm{W}^{*}(A))$ for $A\subseteq\mathcal{M}$. Hence, we restrict our attention to computing the definable and algebraic closures of a $\mathrm{W}^{*}$-subalgebra $\mathcal{A}\subseteq\mathcal{M}$.
@@
Next, we will compute the definable closure of $\mathcal{A}$ in $\mathcal{M}$ whenever $\mathcal{A}\subseteq\mathcal{M}$ is an inclusion of finite-dimensional tracial ***-algebras (as above, $\operatorname{acl}^{\mathcal{M}}(\mathcal{A})=\mathcal{M}$). By Proposition 3.7, $\mathcal{M}$ is $\kappa$-saturated and strongly $\kappa$-homogeneous. Therefore, by Proposition 3.4, $\operatorname{dcl}^{\mathcal{M}}(\mathcal{A})$ is the set of fixed points in $\mathcal{M}$ of the action of automorphisms that fix $\mathcal{A}$ pointwise. The structure theory of finite-dimensional ***-algebras (see for instance [19, §3.1], [42, §3.2]) says there is a decomposition
@@
The next example relates definable closure with normalizers for inclusions of irreducible subfactors $\mathcal{A}\subseteq\mathcal{M}$ that have weak spectral gap, meaning that the only sequences in $\mathcal{M}$ that asymptotically commute with $\mathcal{A}$ are the ones that are asymptotically trivial. Weak spectral gap is an important property in the study of $\mathrm{II}_{1}$ factors that was explicitly formulated in [49], although it was used implicitly earlier through its connection with other rigidity phenomena such as property (T). Recently, Goldbring made several connections between spectral gap and definability of sets [28] which are thematically related to our Proposition 3.12 below. We recall the following characterizations of weak spectral gap; this is in the same spirit as the ultraproduct characterization ([49, Remark 2.2]) of weak spectral gap and the treatment of property Gamma in [23, §3.2.2]. For the reader’s convenience we give a self-contained proof, which is also a good illustration of definition of $\kappa$-saturation.
@@
Many inclusions with weak spectral gap arise from property (T), an important rigidity property defined for groups by Kazhdan [46], for $\mathrm{II}_{1}$ factors by Connes and Jones [14], and for tracial von Neumann algebras by Popa [48]. Indeed, if $\mathcal{A}\subseteq\mathcal{M}$ is an inclusion with $\mathcal{A}^{\prime}\cap\mathcal{M}=\mathbb{C}$ and if $\mathcal{A}$ has property (T), then $\mathcal{A}\subseteq\mathcal{M}$ automatically satisfies the spectral gap condition Lemma 3.11 (1); see e.g. [53, §2.3].
@@
The present author in [39, §4.5] studied the interaction between definable closures and the Jung-Hayes free entropy from [43, 33]. The fact that taking the algebraic closure does not increase the entropy [39, Corollary 4.21] was reminiscent of the behavior of entropy under normalizers (and more generally the singular subspace) obtained earlier by Hayes [33, Theorem 3.1]. Proposition 3.12 gives an explicit connection between these two results by exhibiting some normalizers as examples of algebraic closure. Of course, Hayes’ original results on the singular subspace are far more general than normalizers of subfactors with spectral gap. Nonetheless, this example may provide the first hint as to what the analog of Hayes’ results should be in the context of complete types.
@@ 4. Monge-Kantorovich duality for types and definable predicates
In this section, we study convex definable predicates and prove two of the main results, the non-commutative model-theoretic Monge-Kantorovich duality Theorem 1.1, the analog of Jensen’s inequality Proposition 1.2, and Theorem 1.3 on the displacement interpolation.
@@ 4.1. Proof of duality
As in [24, §3], a key point of the argument for Monge-Kantorovich duality is a non-commutative version of the Legendre transform. As the first step, we prove a weaker version where the predicates $\phi$ and $\psi$ are not required to be convex and only satisfy (1.1) on the domain $D_{\mathbf{r}}$ rather than everywhere.
@@
$\theta_{\epsilon_{1}}^{\mathcal{M}}(\mathbf{y})+\sum_{k=1}^{K}\chi_{k}^{\mathcal{M}}(\mathbf{y})(\theta_{\epsilon_{k+1}}^{\mathcal{M}}(\mathbf{y})-\theta_{\epsilon_{k}}^{\mathcal{M}}(\mathbf{y}))\geq\theta_{\epsilon_{1}}^{\mathcal{M}}(\mathbf{y})+\sum_{n=1}^{N}(\theta_{\epsilon_{k+1}}^{\mathcal{M}}(\mathbf{y})-\theta_{\epsilon_{k}}^{\mathcal{M}}(\mathbf{y}))=\theta_{\epsilon_{K+1}}^{\mathcal{M}}(\mathbf{y})\geq C(\operatorname{tp}^{\mathcal{M}}(\mathbf{y}),\mu).$
@@
$\phi_{0}^{\mathcal{M}}(\mathbf{x})=\sup_{\mathbf{y}\in D_{\mathbf{r}}}\left[\operatorname{Re}\langle\mathbf{x},\mathbf{y}\rangle_{L^{2}(\mathcal{M})^{n}}-\psi_{0}^{\mathcal{M}}(\mathbf{y})\right]$
@@
$\psi_{1}^{\mathcal{M}}(\mathbf{y})=\sup_{\mathbf{x}\in D_{\mathbf{r}}}\left[\operatorname{Re}\langle\mathbf{x},\mathbf{y}\rangle_{L^{2}(\mathcal{M})^{n}}-\phi_{0}^{\mathcal{M}}(\mathbf{x})\right].$
@@
Recall that $\phi_{0}^{\mathcal{M}}(\mathbf{x})$ is the supremum over $\mathbf{z}$ of $\operatorname{Re}\langle\mathbf{x},\mathbf{z}\rangle_{L^{2}(\mathcal{M})^{n}}-\psi_{0}^{\mathcal{M}}(\mathbf{z})$. Since $\lVert\mathbf{z}\rVert_{L^{2}(\mathcal{M})^{n}}\leq|\mathbf{r}|$, each function $\operatorname{Re}\langle\mathbf{x},\mathbf{z}\rangle_{L^{2}(\mathcal{M})^{n}}-\psi_{0}^{\mathcal{M}}(\mathbf{z})$ is $|\mathbf{r}|$-Lipschitz in $\mathbf{x}$. Hence, $\phi_{0}^{\mathcal{M}}$ is $|\mathbf{r}|$-Lipschitz in $\mathbf{x}$ as well. Thus,
@@
$\leq\lVert\widehat{\mathbf{y}}\rVert_{L^{2}(\mathcal{M})^{n}}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{L^{2}(\mathcal{M})^{n}}+\frac{1}{2}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{L^{2}(\mathcal{M})^{n}}^{2}+\frac{1}{2}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{L^{2}(\mathcal{M})^{n}}^{2}$
@@ 4.2. Properties of convex definable predicates
As a corollary, we obtain monotonicity of a convex definable predicate $\phi$ under conditional expectations onto elementary submodels (Proposition 1.2). Monotonicity under conditional expectations in general was a requirement in [24, Definitions 1.2 and 3.6] for $E$-convexity. The condition is motivated by Jensen’s inequality for convex functions on $\mathbb{R}^{n}$ [24, §1.4, Lemma 1.10], as well as the behavior of unitarily invariant convex functions on $M_{n}(\mathbb{C})$ [24, §1.5, Lemma 1.17]. Now let us show that the analogous condition for definable predicates is automatic.
@@ 5.1. Semiconvexity, Semiconcavity, and Differentiability
$\psi_{\epsilon}^{\mathcal{M}}(\mathbf{x},\mathbf{y})=\sup_{\mathbf{z}\in D_{\mathbf{R}}}\frac{(\operatorname{Re}\langle\mathbf{z},\nabla_{\mathbf{x}}\phi(\mathbf{x},\mathbf{y})\rangle_{L^{2}(\mathcal{M})^{n}})^{2}}{\epsilon+\lVert\mathbf{z}\rVert_{L^{2}(\mathcal{M})^{n}}^{2}}$
@@
In Proposition 5.2 (3), we made an additional boundedness assumption that $\nabla_{\mathbf{y}}\phi$ maps a product of domains (operator-norm balls) into a product of domains, which must be verified in applications. We will see examples of definable predicates in the next section where the boundedness condition holds by construction. However, we remark that the boundedness condition is automatic in the case where $\phi$ is a function of $\mathbf{x}$ alone (i.e., there is no $\mathbf{y}$) and $\mathrm{T}\models\mathrm{T}_{\mathrm{II}_{1}}$. This follows from a much more general fact about unitarily equivariant Lipschitz functions on tuples from a $\mathrm{II}_{1}$ factor. A weaker version of this lemma was given earlier in [38] and [37] using a random matrix argument. In the following, $L^{2}(\mathcal{M})_{\operatorname{sa}}$ denotes the real Hilbert space of self-adjoint elements in $L^{2}(\mathcal{M})$. Later on, we will apply this result to $\nabla\phi$ after showing it is Lipschitz.
@@
Let $a$ and $b\in\operatorname{Spec}(F(\mathbf{x}))$, and let $\epsilon>0$. Then because the projection-valued spectral measure of $F(\mathbf{x})$ has support equal to $\operatorname{Spec}(F(\mathbf{x}))$, the spectral projections $1_{[a-\epsilon,a+\epsilon]}(F(\mathbf{x}))$ and $1_{[b-\epsilon,b+\epsilon]}(F(\mathbf{x}))$ are nonzero. Because $\mathcal{M}$ is a $\mathrm{II}_{1}$ factor, there exists a nonzero partial isometry $v$ such that $v^{*}v\leq 1_{[a-\epsilon,a+\epsilon]}(F(\mathbf{x}))$ and $vv^{*}\leq 1_{[b-\epsilon,b+\epsilon]}(F(\mathbf{x}))$ (see e.g. [45, Proposition 6.1.8]). Then
@@
Let $\mathcal{M}$ be a $\mathrm{II}_{1}$ factor and let $F:\mathcal{M}^{n}\to L^{2}(\mathcal{M})^{m}$ be $L$-Lipschitz with respect to $\lVert\cdot\rVert_{L^{2}(\mathcal{M})}$. Let $\mathbf{r}=(r_{1},\dots,r_{n})\in(0,\infty)^{n}$. Let $t=\max_{i}(|\operatorname{tr}^{\mathcal{M}}(F_{i}(0))|)$. Then $F$ maps $D_{\mathbf{r}}^{\mathcal{M}}$ into $(D_{t+9L|\mathbf{r}|}^{\mathcal{M}})^{m}$.

### 09 Mar 2024 v3

@@ 3.1. Definition and Properties
Thus, by Proposition 2.22, $\tilde{\mathcal{P}}\preceq\mathcal{M}$. Next, we claim $\tilde{\mathcal{P}}$ contains $A$. This is because for each domain $D_{0}$, each element of $\mathbf{z}$ in the domain $D_{0}$ is a limit of elements from $\mathbf{w}$ inside $D_{0}^{\mathcal{P}}$, hence the same is true for $\tilde{\mathbf{w}}$ since $d^{\mathcal{M}}(z_{i},w_{j})=d^{\mathcal{M}}(z_{i},\tilde{w}_{j})$ by virtue of the types of $(\mathbf{z},\mathbf{w})$ and $(\mathbf{z},\tilde{\mathbf{w}})$ being the same. Finally, similar reasoning shows that $\tilde{\mathcal{P}}$ does not contain $x$. Indeed, $d(x,D^{\tilde{\mathcal{P}}})$ is the infimum of distances between $x$ and $\tilde{w}_{i}$ for $i\in I^{\prime}$ such that $\tilde{w}_{i}\in D^{\tilde{P}}$. But $d(x,\tilde{w}_{i})=d(y,w_{i})$. Thus, $d(x,D^{\tilde{\mathcal{P}}})=d(y,D^{\mathcal{P}})>0$. Hence, $\tilde{\mathcal{P}}$ is an elementary submodel of $\mathcal{M}$ that contains $A$ but not $x$, which completes the proof. ∎
@@
because $\mathcal{M}\preceq\mathcal{N}_{1}$. Therefore, $\Phi$ is approximately satisfiable in $\mathcal{M}$, and hence $\Phi$ is satisfiable in $\mathcal{M}$ because $\mathcal{M}$ is $\kappa$-saturated. If $x\in D^{\mathcal{M}}$ satisfies $\Phi$, then $\operatorname{tp}^{\mathcal{N}_{1}}(x)=\operatorname{tp}^{\mathcal{M}}(x)=\operatorname{tp}^{\mathcal{N}_{1}}(y)$. Since $\mathcal{N}_{1}$ is strongly $\kappa$-homogeneous, there exists an automorphism $\Theta$ of $\mathcal{N}_{1}$ with $\Theta(x)=y$. By (1) and Proposition 3.3, $x\in\operatorname{acl}^{\mathcal{N}_{1}}(\varnothing)$, which is the intersection of all elementary submodels of $\mathcal{N}_{1}$ by Proposition 3.5. Therefore, because the automorphic image of an elementary submodel is also an elementary submodel, $y=\Theta(x)$ is in the intersection of all elementary submodels of $\mathcal{N}_{1}$, which is $\operatorname{acl}^{\mathcal{N}_{1}}(\varnothing)=\operatorname{acl}^{\mathcal{M}}(\varnothing)=\mathcal{M}$. Therefore, $\mathcal{N}_{1}=\mathcal{M}$ and also $\mathcal{N}=\mathcal{M}$. (3) Let $\mathcal{N}$ be a $\kappa$-saturated and strongly $\kappa$-homogeneous elementary extension of $\mathcal{M}$. Then $\mathcal{M}=\mathcal{N}$ by (2), so we are done. ∎
@@ 3.2. Examples in von Neumann algebras
Definable and algebraic closures have not yet been studied in von Neumann algebras (except for their relationship with $1$-bounded entropy in [40, §4.5]), and therefore, we develop several basic properties and examples here. Given that the definable closure is always a $\mathrm{W}^{*}$-subalgebra by Corollary 3.6, we have $\operatorname{dcl}^{\mathcal{M}}(A)=\operatorname{dcl}^{\mathcal{M}}(\mathrm{W}^{*}(A))$ for $A\subseteq\mathcal{M}$. Hence, we restrict our attention to computing the definable and algebraic closures of a $\mathrm{W}^{*}$-subalgebra $\mathcal{A}\subseteq\mathcal{M}$.
@@
Next, we will compute the definable closure of $\mathcal{A}$ in $\mathcal{M}$ whenever $\mathcal{A}\subseteq\mathcal{M}$ is an inclusion of finite-dimensional tracial ***-algebras (as above, $\operatorname{acl}^{\mathcal{M}}(\mathcal{A})=\mathcal{M}$). By Proposition 3.7, $\mathcal{M}$ is $\kappa$-saturated and strongly $\kappa$-homogeneous. Therefore, by Proposition 3.4, $\operatorname{dcl}^{\mathcal{M}}(\mathcal{A})$ is the set of fixed points in $\mathcal{M}$ of the action of automorphisms that fix $\mathcal{A}$ pointwise. The structure theory of finite-dimensional ***-algebras (see for instance [19, §3.1], [43, §3.2]) says there is a decomposition
@@
The next example relates definable closure with normalizers for inclusions of irreducible subfactors $\mathcal{A}\subseteq\mathcal{M}$ that have weak spectral gap, meaning that the only sequences in $\mathcal{M}$ that asymptotically commute with $\mathcal{A}$ are the ones that are asymptotically trivial. Weak spectral gap is an important property in the study of $\mathrm{II}_{1}$ factors that was explicitly formulated in [50], although it was used implicitly earlier through its connection with other rigidity phenomena such as property (T). Recently, Goldbring made several connections between spectral gap and definability of sets [29] which are thematically related to our Proposition 3.12 below. We recall the following characterizations of weak spectral gap; this is in the same spirit as the ultraproduct characterization ([50, Remark 2.2]) of weak spectral gap and the treatment of property Gamma in [24, §3.2.2]. For the reader’s convenience we give a self-contained proof, which is also a good illustration of definition of $\kappa$-saturation.
@@
Many inclusions with weak spectral gap arise from property (T), an important rigidity property defined for groups by Kazhdan [47], for $\mathrm{II}_{1}$ factors by Connes and Jones [14], and for tracial von Neumann algebras by Popa [49]. Indeed, if $\mathcal{A}\subseteq\mathcal{M}$ is an inclusion with $\mathcal{A}^{\prime}\cap\mathcal{M}=\mathbb{C}$ and if $\mathcal{A}$ has property (T), then $\mathcal{A}\subseteq\mathcal{M}$ automatically satisfies the spectral gap condition Lemma 3.11 (1); see e.g. [54, §2.3].
@@
The present author in [40, §4.5] studied the interaction between definable closures and the Jung-Hayes free entropy from [44, 34]. The fact that taking the algebraic closure does not increase the entropy [40, Corollary 4.21] was reminiscent of the behavior of entropy under normalizers (and more generally the singular subspace) obtained earlier by Hayes [34, Theorem 3.1]. Proposition 3.12 gives an explicit connection between these two results by exhibiting some normalizers as examples of algebraic closure. Of course, Hayes’ original results on the singular subspace are far more general than normalizers of subfactors with spectral gap. Nonetheless, this example may provide the first hint as to what the analog of Hayes’ results should be in the context of complete types.
@@ 4. Monge-Kantorovich duality for types and definable predicates
In this section, we study convex definable predicates and prove three of the main results: the non-commutative model-theoretic Monge-Kantorovich duality Theorem 1.1, the analog of Jensen’s inequality Proposition 1.2, and Theorem 1.3 on the displacement interpolation.
@@ 4.1. Proof of duality
As in [25, §3], a key point of the argument for Monge-Kantorovich duality is a non-commutative version of the Legendre transform. As the first step, we prove a weaker version where the predicates $\phi$ and $\psi$ are not required to be convex and only satisfy (1.1) on the domain $D_{\mathbf{r}}$ rather than everywhere.
@@
$\theta_{\epsilon_{1}}^{\mathcal{M}}(\mathbf{y})+\sum_{k=1}^{K}\chi_{k}^{\mathcal{M}}(\mathbf{y})(\theta_{\epsilon_{k+1}}^{\mathcal{M}}(\mathbf{y})-\theta_{\epsilon_{k}}^{\mathcal{M}}(\mathbf{y}))\geq\theta_{\epsilon_{1}}^{\mathcal{M}}(\mathbf{y})+\sum_{k=1}^{K}(\theta_{\epsilon_{k+1}}^{\mathcal{M}}(\mathbf{y})-\theta_{\epsilon_{k}}^{\mathcal{M}}(\mathbf{y}))=\theta_{\epsilon_{K+1}}^{\mathcal{M}}(\mathbf{y})\geq C(\operatorname{tp}^{\mathcal{M}}(\mathbf{y}),\mu).$
@@
$\phi_{0}^{\mathcal{M}}(\mathbf{x})=\sup_{\mathbf{y}\in D_{\mathbf{r}}^{\mathcal{M}}}\left[\operatorname{Re}\langle\mathbf{x},\mathbf{y}\rangle_{L^{2}(\mathcal{M})^{n}}-\psi_{0}^{\mathcal{M}}(\mathbf{y})\right]$
@@
$\psi_{1}^{\mathcal{M}}(\mathbf{y})=\sup_{\mathbf{x}\in D_{\mathbf{r}}^{\mathcal{M}}}\left[\operatorname{Re}\langle\mathbf{x},\mathbf{y}\rangle_{L^{2}(\mathcal{M})^{n}}-\phi_{0}^{\mathcal{M}}(\mathbf{x})\right].$
@@
Recall that $\phi_{0}^{\mathcal{M}}(\mathbf{x})$ is the supremum over $\mathbf{z}\in D_{\mathbf{r}}^{\mathcal{M}}$ of $\operatorname{Re}\langle\mathbf{x},\mathbf{z}\rangle_{L^{2}(\mathcal{M})^{n}}-\psi_{0}^{\mathcal{M}}(\mathbf{z})$. Since $\lVert\mathbf{z}\rVert_{L^{2}(\mathcal{M})^{n}}\leq|\mathbf{r}|$, each function $\operatorname{Re}\langle\mathbf{x},\mathbf{z}\rangle_{L^{2}(\mathcal{M})^{n}}-\psi_{0}^{\mathcal{M}}(\mathbf{z})$ is $|\mathbf{r}|$-Lipschitz in $\mathbf{x}$. Hence, $\phi_{0}^{\mathcal{M}}$ is $|\mathbf{r}|$-Lipschitz in $\mathbf{x}$ as well. Thus,
@@
$\leq\lVert\widehat{\mathbf{y}}\rVert_{L^{2}(\mathcal{M})^{n}}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{L^{2}(\mathcal{M})^{n}}+\frac{1}{2}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{L^{2}(\mathcal{M})^{n}}^{2}+\frac{1}{2}\lVert\mathbf{y}-\widehat{\mathbf{y}}\rVert_{L^{2}(\mathcal{M})^{n}}^{2}$
@@ 4.2. Properties of convex definable predicates
As a corollary, we obtain monotonicity of a convex definable predicate $\phi$ under conditional expectations onto elementary submodels (Proposition 1.2). Monotonicity under conditional expectations in general was a requirement in [25, Definitions 1.2 and 3.6] for $E$-convexity. The condition is motivated by Jensen’s inequality for convex functions on $\mathbb{R}^{n}$ [25, §1.4, Lemma 1.10], as well as the behavior of unitarily invariant convex functions on $M_{n}(\mathbb{C})$ [25, §1.5, Lemma 1.17]. Now let us show that the analogous condition for definable predicates is automatic.
@@ 5.1. Semiconvexity, Semiconcavity, and Differentiability
$\psi_{\epsilon}^{\mathcal{M}}(\mathbf{x},\mathbf{y})=\sup_{\mathbf{z}\in D_{\mathbf{R}}^{\mathcal{M}}}\frac{(\operatorname{Re}\langle\mathbf{z},\nabla_{\mathbf{x}}\phi(\mathbf{x},\mathbf{y})\rangle_{L^{2}(\mathcal{M})^{n}})^{2}}{\epsilon+\lVert\mathbf{z}\rVert_{L^{2}(\mathcal{M})^{n}}^{2}}$
@@
In Proposition 5.2 (3), we made an additional boundedness assumption that $\nabla_{\mathbf{y}}\phi$ maps a product of domains (operator-norm balls) into a product of domains, which must be verified in applications. We will see examples of definable predicates in the next section where the boundedness condition holds by construction. However, we remark that the boundedness condition is automatic in the case where $\phi$ is a function of $\mathbf{x}$ alone (i.e., there is no $\mathbf{y}$) and $\mathrm{T}\models\mathrm{T}_{\mathrm{II}_{1}}$. This follows from a much more general fact about unitarily equivariant Lipschitz functions on tuples from a $\mathrm{II}_{1}$ factor. A weaker version of this lemma was given earlier in [38, Lemma 11.5.4] using a random matrix argument. In the following, $L^{2}(\mathcal{M})_{\operatorname{sa}}$ denotes the real Hilbert space of self-adjoint elements in $L^{2}(\mathcal{M})$. Later on, we will apply this result to $\nabla\phi$ after showing it is Lipschitz.
@@
Let $a$ and $b\in\operatorname{Spec}(F(\mathbf{x}))$, and let $\epsilon>0$. Then because the projection-valued spectral measure of $F(\mathbf{x})$ has support equal to $\operatorname{Spec}(F(\mathbf{x}))$, the spectral projections $1_{[a-\epsilon,a+\epsilon]}(F(\mathbf{x}))$ and $1_{[b-\epsilon,b+\epsilon]}(F(\mathbf{x}))$ are nonzero. Because $\mathcal{M}$ is a $\mathrm{II}_{1}$ factor, there exists a nonzero partial isometry $v$ such that $v^{*}v\leq 1_{[a-\epsilon,a+\epsilon]}(F(\mathbf{x}))$ and $vv^{*}\leq 1_{[b-\epsilon,b+\epsilon]}(F(\mathbf{x}))$ (see e.g. [46, Proposition 6.1.8]). Then
@@
Let $\mathcal{M}$ be a $\mathrm{II}_{1}$ factor and let $F:\mathcal{M}^{n}\to L^{2}(\mathcal{M})^{m}$ be $L$-Lipschitz with respect to $\lVert\cdot\rVert_{L^{2}(\mathcal{M})}$ and equivariant under unitary conjugation. Let $\mathbf{r}=(r_{1},\dots,r_{n})\in(0,\infty)^{n}$. Let $t=\max_{i}(|\operatorname{tr}^{\mathcal{M}}(F_{i}(0))|)$. Then $F$ maps $D_{\mathbf{r}}^{\mathcal{M}}$ into $(D_{t+9L|\mathbf{r}|}^{\mathcal{M}})^{m}$.